### v1

# Negative Yields Positive: Unified Dual-Path Adapter for Vision-Language Models

Abstract
Recently, large-scale pre-trained Vision-Language Models (VLMs) have demonstrated great potential in learning open-world visual representations, and exhibit remarkable performance across a wide range of downstream tasks through efficient fine-tuning. In this work, we innovatively introduce the concept of dual learning into fine-tuning VLMs, i.e., we not only learn what an image is, but also what an image isn’t. Building on this concept, we introduce a novel DualAdapter approach to enable dual-path adaptation of VLMs from both positive and negative perspectives with only limited annotated samples. In the inference stage, our DualAdapter performs unified predictions by simultaneously conducting complementary positive selection and negative exclusion across target classes, thereby enhancing the overall recognition accuracy of VLMs in downstream tasks. Our extensive experimental results across 15 datasets validate that the proposed DualAdapter outperforms existing state-of-the-art methods on both few-shot learning and domain generalization tasks while achieving competitive computational efficiency. Code is available at https://github.com/zhangce01/DualAdapter.
Keywords:
Vision-Language Models CLIP Few-Shot Classification Domain Generalization Transfer Learning
1 Introduction
Large-scale pre-trained Vision-Language Models (VLMs), such as CLIP [48] and ALIGN [24], provide a new paradigm for multi-modal learning and open-world visual recognition [78, 8]. Recently, researchers have demonstrated the robustness and effectiveness of these VLMs in interpreting complex visual and textual inputs, as evidenced by their superior performance in a variety of vision-language tasks, e.g., image captioning [73, 21, 60], visual question answering [52, 25, 31], and other domains [7, 85, 54, 58].
Current VLMs excel in zero-shot inference [86, 15], but struggle in transferring to specific downstream tasks, primarily due to domain shifts and mismatches between pre-training and task-specific data [10, 42]. Additionally, the extensive number of parameters and demanding computational costs make it intractable for individual practitioners to re-train or fine-tune the entire models to enhance their task-specific performance [84, 12]. To address this, recent studies have focused on designing efficient fine-tuning strategies that maintain the core pre-trained encoders frozen while incorporating lightweight adapter [12, 79] modules or optimizing prompts[84, 83]. These methods enable effective task-specific knowledge extraction from a limited amount of annotated training samples, on top of the inherited foundational knowledge from pre-training.
Figure 1: Dual-path inference examples. We show four sample images from diverse recognition tasks, along with their similarities (scaled by 100) to positive/negative text descriptions after few-shot dual-path adaptation. Intuitively, an image is more likely to be classified into a specific class if it exhibits higher positive similarity and lower negative similarity. We can observe that while the positive branch alone may struggle to distinguish among some closely related classes, incorporating the negative path, which eliminates certain incorrect classes, enhances its ability to accurately identify the true class.
In this work, we introduce the innovative concept of dual inference in the context of fine-tuning VLMs. Take, for instance, a basic binary pet classification task: distinguishing cats from dogs. To identify an image as a cat, we can approach the task from both the positive and negative perspectives: one might directly identify the image as a cat (selection), or alternatively, one can also first determine that it is not a dog (exclusion), thereby indirectly concluding it’s a cat. These approaches, namely, positive selection and negative exclusion, complement each other and collectively enhance the overall classification accuracy in general classification tasks.
Building on this concept, we introduce a novel DualAdapter approach for few-shot adaptation of VLMs that conducts positive selection and negative exclusion across a set of target classes. While CLIP typically performs classification tasks using prompts such as “A photo of a {CLASS}”, we recognize its ability to understand negation, employing prompts like “A photo of no {CLASS}”. Utilizing this capability, we design our DualAdapter to adapt VLMs from a dual perspective: we propose to learn not only what an image is but also what an image isn’t by matching the input image feature with the positive and negative text features. Intuitively, our approach first determines a set of likely candidate classes through image classification but, in a second step, refines these classification logits by asking the inverse question of what an image isn’t across the entire set of target classes. In Fig. 1, we showcase several dual-path inference examples to illustrate that when positive selection struggles with distinguishing between similar classes, employing negative exclusion can provide further insights to improve recognition accuracy.
Moreover, in DualAdapter, we symmetrically apply dual-path adaptation to the visual modality, treating image features of a specific class as positive and those from all other classes as negative. Recognizing the variability in image quality and representativeness, where some image features are noisy and not equally representative of their respective classes, we implement an unsupervised similarity-based label refinement technique. This method assigns confidence scores to each image, reducing the impact of outliers or less representative samples during few-shot adaptation, ensuring more accurate and reliable classification.
To empirically validate DualAdapter, we perform thorough evaluations on few-shot learning and domain generalization tasks using 15 diverse recognition datasets. These experimental results demonstrate the effectiveness of DualAdapter in adapting VLMs for downstream tasks and verify its superior robustness to distribution shifts. Additionally, our efficiency benchmarks indicate that DualAdapter achieves competitive efficiency compared to other state-of-the-art methods.
In summary, our primary contributions are listed as follows:
  • $\bullet$
    We explore and exploit the negative inference capabilities of VLMs, and introduce a unique dual-path adaptation approach (i.e., DualAdapter) for CLIP.
  • $\bullet$
    To mitigate the impact of noisy samples in few-shot adaptation, we introduce a novel unsupervised similarity-based label refinement technique.
  • $\bullet$
    Through extensive experiments, we demonstrate that our DualAdapter outperforms other state-of-the-art methods in both adaptation performance and generalizability across 15 diverse recognition datasets.
2 Related Work
Vision-Language Models. In recent years, extensive Vision-Language Models (VLMs) have been developed to bridge the vision and language modalities through large-scale pre-training [48, 24]. Generally, these pre-trained VLMs can be classified into four categories based on their pre-training objectives: masked language modeling [37, 30, 27], masked region prediction [37, 30, 59], image-text matching [30, 27, 61, 31], and contrastive learning [48, 24, 74, 77, 33]. Researchers have demonstrated that these pre-trained VLMs exhibit remarkable open-world visual understanding capabilities [8], achieving state-of-the-art performance across a range of open-world tasks, such as zero-shot learning [48, 51, 41], open-world segmentation [34, 72, 49], and open-world detection [53, 16, 9]. In this work, we mainly focus on contrastive-based VLMs, especially CLIP [48], which has become the current mainstream approach in this field. Nonetheless, we believe that the core concept of this work is adaptable and can be extended to other VLMs as well.
Efficient Fine-Tuning for VLMs. Given the substantial size of the VLMs, recent research efforts [84, 12, 79, 80] are focusing on the development of lightweight fine-tuning techniques to efficiently adapt VLMs for downstream visual tasks. These methods can generally be divided into two primary categories: prompt-based learning and adapter-style fine-tuning. Specifically, prompt-based learning methods [84, 83, 2, 62] aim to optimize the input prompts from downstream data, while adapter-style fine-tuning approaches [12, 79, 64, 75, 88] directly tune the extracted visual and textual representations. Our proposed DualAdapter follows the line of adapter-style fine-tuning, but it distinguishes itself from previous work [79, 75] by incorporating negative adapters for the first time. Moreover, we empirically demonstrate that leveraging negative cues from CLIP can effectively improve both few-shot classification performance and generalization capability.
Negative Learning. Given that obtaining typical positive labels (which indicate the categories an image belongs to) can be costly and labor-intensive, researchers have proposed an alternative approach known as the indirect negative learning [23, 76] paradigm. This method focuses on learning from easily accessible negative/complementary labels, which specify the categories to which an image does not belong. In recent years, negative learning has been effectively applied to various vision applications, such as image recognition [23, 28, 13], few-shot learning [22, 69], and semantic segmentation [67, 47]. In this work, empowered by the strong image-text association capabilities of the advanced pre-trained CLIP [48] model, we enable negative learning within CLIP without the need for explicit negative labels. Instead, we design negative prompts and negative image samples to guide the CLIP model to predict from a negative perspective, which in turn enhances its overall positive prediction capability. To the best of our knowledge, there are only two concurrent approaches (i.e., CLIPN [66] and LSN [43]) that work in a similar manner: they both enable the negative predictions of CLIP to identify unknown out-of-distribution samples. However, we formulate negative learning from a different and more fundamental perspective, i.e., using negative prediction to directly improve the accuracy of positive prediction.
Contrastive Learning. We also recognize the underlying concept behind our proposed DualAdapter has parallels with existing supervised contrastive learning literature [14, 82, 26]: Our method also aims to maximize similarities between positive image-text/image-image pairs, while simultaneously minimizing similarities between negative pairs during few-shot adaptation. However, our work focuses on a fundamentally different dual learning perspective, i.e., we explicitly train two distinct positive and negative classifiers for ensembling using a single image input at a time, rather than applying supervised contrastive loss to train a single classifier with image pairs as inputs. It is also important to note that contrastive learning is not directly applicable in our fine-tuning setting because we keep the CLIP image encoder frozen. Thereby, the extracted image features are fixed and cannot be pulled together or pushed apart.
Figure 2: An overview of our proposed DualAdapter. We introduce four adapters, each of which builds a cache with positive/negative text/image features that are updated during few-shot training. Given an image to be classified, the classification logit for a specific class increases when the image feature closely aligns with the corresponding features in the positive cache and diverges from those in the negative cache.
3 DualAdapter: Unified Dual-Path Adaptation of VLMs
We introduce a novel framework DualAdapter, as illustrated in Fig. 2, to enable efficient dual-path adaptation of VLMs. Our DualAdapter consists of four adapters: two that positively adapt CLIP to enhance its ability to accurately identify the true class of an input image, and two that negatively adapt CLIP to effectively exclude incorrect candidate classes. Furthermore, we formulate few-shot learning as a noisy problem, and propose an unsupervised similarity-based label refinement approach to select high-quality representative samples to more effectively adapt CLIP to downstream tasks. In the following sections, we first briefly review zero-shot CLIP, then introduce each component of our DualAdapter in detail.
3.1 Preliminary: A Revisit of CLIP
By leveraging contrastive training on billions of image-text pairs from the Internet, CLIP has shown an outstanding ability to align vision and language modalities [48]. In this work, we employ CLIP’s pre-trained visual encoder $\mathcal{F}_{\mathsf{V}}$ and textual encoder $\mathcal{F}_{\mathsf{T}}$ to map the images and textual descriptions into a shared $d$-dimensional embedding space. For a $C$-class classification task, CLIP performs zero-shot predictions by assessing the similarity between the image feature and the text feature specific to each class $c\in\{1,...,C\}$ as follows:
$f_{\mathcal{I}}=\mathcal{F}_{\mathsf{V}}(\mathcal{I}),\quad f_{t_{c}}=\mathcal{F}_{\mathsf{T}}(\mathcal{T}_{c}),\quad\mathbb{P}(y=c|\mathcal{I})=\frac{\exp\left(\mathrm{cos}\left(f_{t_{c}},f_{\mathcal{I}}\right)/t\right)}{\sum\nolimits_{t^{\prime}}{\exp\left(\mathrm{cos}\left(f_{t^{\prime}},f_{\mathcal{I}}\right)/t\right)}},$ (1)
[…]
To streamline this process, a weight matrix can be precomputed and stored in a textual cache, which concatenates the textual features associated with each class, denoted as $\mathsf{T}_{\mathtt{cache}}=[f_{t_{1}}\,f_{t_{2}}\,\cdots\,f_{t_{C}}]^{\top}\in\mathbb{R}^{C\times d}$. Subsequently, we can efficiently obtain the logit $\mathcal{S}$ and the final prediction $\mathbb{P}(y|\mathcal{I})$ via vectorized computation:
$\mathcal{S}=f_{\mathcal{I}}\mathsf{T}_{\mathtt{cache}}^{\quad\,\,\,\,\top}\in\mathbb{R}^{1\times C},\quad\mathbb{P}(y|\mathcal{I})=\mathsf{Softmax}\left(\mathcal{S}\right).$ (2)
3.2 Positive Adaptation
Although the zero-shot CLIP model is trained on web-scale datasets and encodes extensive prior knowledge, it cannot be effectively applied to downstream tasks due to the presence of domain shifts and a deficiency in task-specific knowledge. To address this, we efficiently leverage few-shot annotated samples to distill new, task-specific knowledge. As we discussed in Section 1, our DualAdapter enables few-shot adaptation from both positive and negative perspectives. In this section, we first focus on the positive side, i.e., refining CLIP’s ability to more precisely predict the true class of the input image in a specific downstream task.
Positive Textual Adapter. As shown in Eq. (2), CLIP can make zero-shot predictions utilizing a textual cache, which stores the textual features from positive text descriptions (e.g., “A photo of a {CLASS}”). To avoid ambiguity, we denote this cache as $\mathsf{T}_{\mathtt{cache}}^{+}=[f_{t_{1}}^{+}\,f_{t_{2}}^{+}\,\cdots\,f_{t_{C}}^{+}]^{\top}\in\mathbb{R}^{C\times d}$. With a small set of annotated training images, we can improve this cache-based classifier by enhancing the inter-modal alignment between textual features and the few-shot image features from the same class. Specifically, we follow TaskRes [75] to achieve this by introducing a group of learnable parameters $\mathcal{R}_{\mathsf{T}}^{+}\in\mathbb{R}^{C\times d}$. By incorporating these parameters with the textual cache $\mathsf{T}_{\mathtt{cache}}^{+}$ via element-wise addition in a residual approach, it integrates task-specific knowledge and enhances the inter-modal classifier:
$\mathsf{T}_{\mathtt{cache}}^{+}\leftarrow\mathsf{Normalize}\left(\mathsf{T}_{\mathtt{cache}}^{+}+\mathcal{R}_{\mathsf{T}}^{+}\right),\quad\mathcal{S}_{\mathsf{T}}^{+}=f_{\mathcal{I}}\mathsf{T}_{\mathtt{cache}}^{+\quad\top}\in\mathbb{R}^{1\times C}.$ (3)
Here, $\mathsf{Normalize}$ denotes the L2-normalization applied to each row of the matrix.
Positive Visual Adapter. Inspired by Tip-Adapter [79], we further extend the recognition capability of the CLIP model by constructing a visual cache-based classifier, which operates by measuring intra-modal similarities between the input image feature and few-shot image features. Given a $C$-class $K$-shot training dataset from a specific downstream task, we store all $CK$ image features in a precomputed visual cache, denoted as $\mathsf{V}_{\mathtt{cache}}^{+}=[f_{v_{1}}^{(1)}\,f_{v_{1}}^{(2)}\,\cdots\,f_{v_{C}}^{(K)}]^{\top}\in\mathbb{R}^{CK\times d}$. Their one-hot labels are also correspondingly recorded in $L\in\mathbb{R}^{CK\times C}$. Given an image feature $f_{\mathcal{I}}$ to be classified, we calculate its image-image affinities $\mathcal{A}^{+}$ with all the few-shot training images, then multiplied by their corresponding one-hot labels $L$ to obtain the classification logit $\mathcal{S}_{\mathsf{V}}^{+}$:
$\mathcal{A}^{+}=\exp\left(-\beta\left(1-f_{\mathcal{I}}\mathsf{V}_{\mathtt{cache}}^{+\quad\top}\right)\right)\in\mathbb{R}^{1\times CK},\quad\mathcal{S}_{\mathsf{V}}^{+}=\alpha\mathcal{A}^{+}L\in\mathbb{R}^{1\times C},$ (4)
where $\alpha$ represents a balance factor and $\beta$ denotes a modulating hyper-parameter. Empirically, the value of $\alpha$ should be increased in cases of larger domain shifts, where more domain-specific few-shot knowledge is needed for adaptation.
Symmetrically, we also introduce a set of learnable parameters $\mathcal{R}_{\mathsf{V}}^{+}\in\mathbb{R}^{C\times d}$, which are broadcast to $\mathbb{R}^{CK\times d}$ and added to the positive visual cache:
$\mathsf{V}_{\mathtt{cache}}^{+}\leftarrow\mathsf{Normalize}\left(\mathsf{V}_{\mathtt{cache}}^{+}+\mathcal{R}_{\mathsf{V}}^{+}\right).$ (5)
3.3 Negative Adaptation
Having introduced the positive branch of our DualAdapter, we now explore the negative inference capabilities of VLMs, adapting them to downstream tasks from a negative perspective. Specifically, we craft negative prompts and generate negative image prototypes to direct CLIP towards more confidently excluding incorrect classes based on the given input image.
Negative Textual Adapter. Recall that the positive textual cache, denoted as $\mathsf{T}_{\mathtt{cache}}^{+}$, is constructed based on the positive class descriptor prompt such as “A photo of a {CLASS}.” Conversely, we introduce a set of negative prompts (e.g., “A photo of no {CLASS},” “A photo without {CLASS}”) that convey meanings opposite in semantics [^1] The comprehensive collection of positive and negative prompts utilized for each dataset is available in the Supplementary Materials.. By leveraging the textual features $f_{t}^{-}$ derived from the negative text descriptions linked with these prompts, we again precompute a weight matrix and construct negative textual cache $\mathsf{T}_{\mathtt{cache}}^{-}=[f_{t_{1}}^{-}\,f_{t_{2}}^{-}\,\cdots\,f_{t_{C}}^{-}]^{\top}\in\mathbb{R}^{C\times d}$. Intuitively, if the feature of an input image $f_{\mathcal{I}}$ closely resembles the negative textual feature for a specific class $c$, it strongly suggests that the image does not belong to class $c$, i.e., $\mathbb{P}(y=c|\mathcal{I})\propto 1-f_{\mathcal{I}}^{\top}f_{t_{c}}^{-}$. Considering the zero-shot CLIP model is not designed to perform this kind of negative prediction, we incorporate a learnable residual $\mathcal{R}_{\mathsf{T}}^{-}\in\mathbb{R}^{C\times d}$ to refine the negative text embeddings through supervised task-specific training. The prediction logit is subsequently given by:
$\mathsf{T}_{\mathtt{cache}}^{-}\leftarrow\mathsf{Normalize}\left(\mathsf{T}_{\mathtt{cache}}^{-}+\mathcal{R}_{\mathsf{T}}^{-}\right),\quad\mathcal{S}_{\mathsf{T}}^{-}=\delta_{\mathsf{T}}\left(\mathbf{1}-f_{\mathcal{I}}\mathsf{T}_{\mathtt{cache}}^{-\quad\top}\right)\in\mathbb{R}^{1\times C},$ (6)
where $\delta_{\mathsf{T}}$ is a fixed scaling parameter that adjusts $\mathcal{S}_{\mathsf{T}}^{-}$ to match the mean value of $\mathcal{S}_{\mathsf{T}}^{+}$, and $\mathbf{1}\in\mathbb{R}^{1\times C}$ denotes an all-ones $C$-dimensional vector.
Negative Visual Adapter. In the visual domain, we also similarly pursue some negative image features to enable negative predictions. To achieve this, we naturally treat all $(C-1)K$ images from the other $C-1$ classes as negative samples for a given class $c$. These negative image samples hold the same property as the negative text descriptions: they are explicitly selected to be non-representative of a specific class, which means that the higher the similarity to them, the lower the probability that the image is classified to that class. To mitigate individual biases, we randomly select one image from each of the $C-1$ classes and compute the average of their extracted features to represent the negative image prototypes [56]. In this way, we can get a total of $K$ negative image prototypes for each of the $C$ classes, thereby constructing a negative visual cache $\mathsf{V}_{\mathtt{cache}}^{-}\in\mathbb{R}^{CK\times d}$. Similar to Eq. (4), we compute the affinities between the input image and those negative image prototypes, and calculate the classification logit:
$\mathcal{A}^{-}\!=\!\delta_{\mathsf{V}}\exp\left(-\beta\left(\mathbf{1}-\left(\mathbf{1}-f_{\mathcal{I}}\mathsf{V}_{\mathtt{cache}}^{-\quad\top}\right)\right)\right)$ $\!=\!\delta_{\mathsf{V}}\exp\left(-\beta f_{\mathcal{I}}\mathsf{V}_{\mathtt{cache}}^{-\quad\top}\right)\in\mathbb{R}^{1\times CK},$ (7)
$\mathcal{S}_{\mathsf{V}}^{-}=\alpha\mathcal{A}^{-}L$ $\in\mathbb{R}^{1\times C}.$ (8)
where $\delta_{\mathsf{V}}$ is another fixed scaling parameter that adjusts $\mathcal{A}^{-}$/$\mathcal{S}_{\mathsf{V}}^{-}$ to match the mean value of $\mathcal{A}^{+}$/$\mathcal{S}_{\mathsf{V}}^{+}$, $L\in\mathbb{R}^{CK\times C}$ denotes the corresponding one-hot labels.
To further encode the diverse fine-grained visual knowledge associated with the specific downstream task, we introduce an extra set of learnable residual parameters, $\mathcal{R}_{\mathsf{V}}^{-}\in\mathbb{R}^{C\times d}$, which is similarly broadcast to refine the negative visual cache:
$\mathsf{V}_{\mathtt{cache}}^{-}\leftarrow\mathsf{Normalize}\left(\mathsf{V}_{\mathtt{cache}}^{-}+\mathcal{R}_{\mathsf{V}}^{-}\right).$ (9)
3.4 Unified Dual-Path Adapter
To effectively adapt VLMs to downstream tasks, our DualAdapter combines the functionalities of all four introduced adapters. As illustrated in Figure 3 (Left), we provide an overview of the specific example features stored in the cache for the class dog. By measuring the similarities between the input image feature and these cached example features, our DualAdapter produces a unified final prediction of the probability that the input image represents a dog. The same process is consistently applied across all classes to perform recognition tasks.
DualAdapter Inference. To derive the final classification scores, we aggregate the four outputs from dual-path (i.e., positive and negative) adapters spanning both textual and visual modalities, expressed as:
$\mathcal{S}_{\mathsf{final}}=\lambda\left(\mathcal{S}_{\mathsf{T}}^{+}+\mathcal{S}_{\mathsf{V}}^{+}\right)+(1-\lambda)\left(\mathcal{S}_{\mathsf{T}}^{-}+\mathcal{S}_{\mathsf{V}}^{-}\right).$ (10)
Here, $\lambda$ serves as a tuning hyper-parameter to balance the contribution of positive and negative adapter logits. Throughout the training process, the collection of learnable parameters =$\mathcal{R}=\{\mathcal{R}_{\mathsf{T}}^{+},\mathcal{R}_{\mathsf{V}}^{+},\mathcal{R}_{\mathsf{T}}^{-},\mathcal{R}_{\mathsf{V}}^{-}\}$ is updated via stochastic gradient descent, employing a cross-entropy loss function for optimization.
Figure 3: Visualization of cosine similarities on the validation set of ImageNet [4]. (Left) Example features of the class dog stored in the cache; (Middle) Distribution of pairwise image-text similarities with dual-path features in textual adapters; (Right) Distribution of pairwise image-image similarities with dual-path features in visual adapters.
Effectiveness of Four Adapters. Fig. 3 qualitatively shows the effectiveness of the four adapters we designed. In the middle and right sub-figure, we visualize the distributions of the pairwise cosine similarities between the features of the input image and the textual/visual features stored in the adapters on ImageNet [4] validation set. From Fig. 3, we can observe the following statistical patterns: (1) As we expected, the input image feature is more similar with positive features from the same class (blue $>$ orange) and negative features from other classes (red $>$ green) across both two modalities; (2) Within the textual modality, negative features tend to be less similar to the input image compared to positive features (red $<$ blue, green $<$ orange), since the negative prompts are less common in CLIP [48] training corpus; (3) Within the visual modality, the similarity distribution of negative features occupies an intermediate position between the positive features of the same and different classes (orange $<$ green/red $<$ blue). This is because the negative visual features are constructed by averaging the image features across all other classes, inherently leading to a more generic representation that lacks the distinctiveness characteristic of positive features within a single class.
3.5 Similarity-Based Label Refinement
Previous adapter-style fine-tuning methods [79, 64] generally assume that the few-shot training samples are meticulously curated to provide accurate representation for each class. However, in practical scenarios, this assumption often does not hold true. We recognize that not every image is of high quality or equally representative of its respective class, which can be particularly detrimental in the context of few-shot learning, where the availability of training images is extremely limited. For example, we provide the t-SNE [38] visualization of few-shot image features from the OxfordPets [45] dataset in Fig. 4 (Left). We argue that the few-shot samples are highly noisy: there is even a data point from the Basset Hound class that overlaps with the prototype (i.e., average feature) of Cocker Spaniel.
In response to this, we devise an unsupervised similarity-based label refinement process to address this challenge. Recall that in our visual adapter, we calculate the image-image affinities between the input image and few-shot training images, with all few-shot images being assigned equal weights [^2] Each few-shot image is assigned a one-hot label, corresponding to a weight of 1. as detailed in Eq. (4). As we discussed, this approach is possibly problematic in the noisy few-shot learning setting. Consequently, we extend the visual adapter to assign non-uniform confidence scores according to pairwise similarities, effectively downweighting outliers and emphasizing more representative samples.
Specifically, our refinement process is based on an intuitive assumption: the representative image feature is closer to other image features from the same class than those low-quality outliers [26]. Based on this assumption, given the $K$-shot image features $\{f_{v_{c}}^{(i)}\}_{i=1}^{K}$ from a specific class $c$, we calculate the average cosine similarities of each image feature to others:
$\mathsf{d}_{c}^{(i)}=\frac{1}{K-1}\sum_{j,\,j\neq i}\cos\left(f_{v_{c}}^{(i)},f_{v_{c}}^{(j)}\right)=\frac{1}{K-1}\sum_{j,\,j\neq i}f_{v_{c}}^{(i)\top}f_{v_{c}}^{(j)}.$ (11)
We then assign non-uniform confidence scores $w_{c}^{(i)}$ and compute refined label values $\ell_{c}^{(i)}$ for all $K$-shot image features based on their similarities to others:
$w_{c}^{(i)}=\frac{\exp(\mathsf{d}_{c}^{(i)}/\tau)}{\sum_{i^{\prime}}\exp(\mathsf{d}_{c}^{(i^{\prime})}/\tau)},\quad\ell_{c}^{(i)}=Kw_{c}^{(i)},$ (12)
where $\tau$ is a temperature hyper-parameter to control the intensity of our refinement. In particular, as $\tau\rightarrow\infty$, our refinement becomes inactive, assigning a label value of 1 to all instances. Conversely, as $\tau\rightarrow 0$, our refinement process becomes more selective, assigning the highest label value exclusively to the most representative image feature, while all other features are assigned 0.
This refinement process is applied across each class, wherein the original one-hot labels $L$, are weighted by the new label values computed in Eq. (12), resulting in the updated labels $\mathbb{L}$. This refinement process is incorporated into both the positive and negative visual adapters; therefore we can rewrite Eqs. (4) and (8) as:
$\mathcal{S}_{\mathsf{V}}^{+}=\alpha\mathcal{A}^{+}\mathbb{L}\in\mathbb{R}^{1\times C},\quad\mathcal{S}_{\mathsf{V}}^{-}=\alpha\mathcal{A}^{-}\mathbb{L}\in\mathbb{R}^{1\times C}.$ (13)
Figure 4: Motivation for the proposed similarity-based label refinement. (Left) t-SNE [38] visualization of visual features for 4 representative classes from the OxfordPets [45] dataset, with prototype [56] of each class denoted by the crossmark $\times$; (Middle) Our refined labels for 4 example images from class Beagle; (Right) An illustrative plot of the modeled distributions using one-hot labels and our refined labels for the Beagle class.
Effectiveness of Label Refinement. In Fig. 4 (Middle), we present our refined label values for 4 example images from class Beagle in OxfordPets [45] dataset. Notably, our unsupervised refinement approach functions by measuring the pairwise similarities among 8-shot images from class Beagle, without requiring any extra parameters or additional training. Our similarity-based refinement process allocates higher values to the more representative samples at the top, and lower values to the lower-quality images at the bottom, where the Beagle is either obscured or presented in black and white. In Fig. 4 (Right), we provide an illustration of the modeled visual feature distribution for the Beagle class using one-hot labels and our refined labels. Through our refinement process, we are better equipped to capture the underlying feature distribution (typically, a normal distribution [68, 81]) of the Beagle class, whereas the use of one-hot labels restricts us to modeling merely a uniform distribution.
4 Experiments
In this section, we conduct extensive experiments on two tasks across 15 datasets. These results demonstrate that our DualAdapter outperforms other state-of-the-art methods, exhibiting enhanced adaptation and generalization capabilities.
4.1 Experimental Settings
Tasks and Datasets. To validate the effectiveness of our proposed DualAdapter, we evaluate our proposed method on two standard benchmarking tasks: few-shot learning and domain generalization, respectively. For few-shot learning task, we comprehensively evaluate our method on 11 well-known image classification benchmarks: ImageNet [4], Caltech101 [11], OxfordPets [45], StandfordCars [29], Flowers102 [44], Food101 [1], FGVCAircraft [39], DTD [3], SUN397 [71], EuroSAT [18], and UCF101 [57]. For domain generalization, we investigate the generalization capability of our DualAdapter on 4 variant datasets of ImageNet: ImageNet-V2 [50], ImageNet-Sketch [65], ImageNet-A [20], and ImageNet-R [19].
Implementation Details. Following previous works [83, 79], we adopt ResNet-50 [17] backbone as the visual encoder of CLIP in our experiments by default. We adopt prompt ensembling, leveraging textual prompts from both CLIP [48] and CuPL [46] to enhance model performance. For the negative prompts we used for each dataset, please kindly refer to the Supplementary Materials. We set the hyper-parameters $\lambda$ and $\tau$ as $0.75$ and $1$, respectively. Our DualAdapter is trained using the AdamW [36] optimizer with a cosine scheduler [35]. The batch size is set to 256. For $\mathcal{R}_{\mathsf{T}}^{+}$ and $\mathcal{R}_{\mathsf{V}}^{+}$, the learning rate is set to $0.0001$, while for $\mathcal{R}_{\mathsf{T}}^{-}$ and $\mathcal{R}_{\mathsf{V}}^{-}$, the learning rate is set to $0.0005$. Additionally, our model is trained for 200 epochs on the EuroSAT [18] dataset, and for 20 epochs on all other datasets. To ensure the reliability of our results, we perform each experiment three times using different initialization seeds and report the mean accuracy achieved. All experiments are conducted on a single NVIDIA RTX 6000 Ada GPU.
Baselines. We compare our proposed method with the following state-of-the-art methods: zero-shot and linear probe CLIP [48], CoOp [84], CoCoOp [83], ProGrad [87], CLIP-Adapter [12], Tip-Adapter-F [79], TPT [55], TaskRes [75], and GraphAdapter [32]. For a fair comparison, we directly report the results of these baselines from their respective original papers.
4.2 Few-Shot Learning
In Fig. 5, we compare the few-shot learning performance of our DualAdapter with other state-of-the-art methods on 11 image classification datasets. In the top-left sub-figure, we also present the average classification accuracy across all 11 datasets. The results indicate that our DualAdapter consistently outperforms other methods across various few-shot learning settings by substantial margins. Moreover, our DualAdapter demonstrates more pronounced performance improvements in specialized classification tasks, such as satellite image and texture classification on the EuroSAT [18] and DTD [3] datasets. With 16-shot training on these two datasets, our DualAdapter surpasses Tip-Adapter-F [79] by a notable 3.56% and 3.50%, and enhances the performance of zero-shot CLIP [48] significantly by 45.56% and 30.50%, respectively. For full numerical results, please refer to our Supplementary Materials. Overall, the consistently superior performance on 11 datasets fully demonstrates the effectiveness of our DualAdapter.
Figure 5: Performance comparisons on few-shot learning on 11 image classification datasets. For each dataset, we report the mean accuracy and 95% confidence interval over 3 random seeds of our DualAdapter on 1-/2-/4-/8-/16-shot settings.
4.3 Robustness to Natural Distribution Shifts
Table 1: Performance comparison on robustness to distribution shifts. All the models are trained on 16-shot ImageNet [4] and directed tested on the OOD target datasets. The best results are in bold and the second best are underlined.
Method Source Target
ImageNet -V2 -Sketch -A -R Avg.
Zero-Shot CLIP [48] 60.33 53.27 35.44 21.65 56.00 41.59
Linear Probe CLIP [48] 56.13 45.61 19.13 12.74 34.86 28.09
CoOp [84] 62.95 55.40 34.67 23.06 56.60 42.43
CoCoOp [83] 62.71 55.72 34.48 23.32 57.74 42.82
ProGrad [87] 62.17 54.70 34.40 23.05 56.77 42.23
TPT [55] 60.74 54.70 35.09 26.67 59.11 43.89
TaskRes [75] 64.75 56.47 35.83 22.80 60.70 43.95
GraphAdapter [32] 64.94 56.58 35.89 23.07 60.86 44.10
DualAdapter (Ours) 66.52 57.87 36.38 25.73 61.12 45.28
In Table 1, we compare the generalization performance of our DualAdapter with other state-of-the-art methods in the presence of distribution shifts. Specifically, all the models are trained solely on 16-shot ImageNet [4], and directly tested on 4 out-of-distribution ImageNet variant datasets. As shown in Table 1, our DualAdapter not only achieves state-of-the-art performance on the source dataset but also attains an average performance gain of 1.18% across 4 out-of-distribution (OOD) target datasets. These experimental results demonstrate that by enabling adaptation through both positive and negative paths, our DualAdapter can more effectively prevent overfitting to the source dataset, thereby enhancing robustness against distribution shifts.
4.4 Ablation Studies
[…]
Table 2: Ablation studies for different variants of DualAdapter. We evaluate the few-shot adaptation capabilities of four DualAdapter variants on ImageNet [4].
# Method $\mathcal{R}_{\mathsf{T}}^{+}$ $\mathcal{R}_{\mathsf{T}}^{-}$ $\mathcal{R}_{\mathsf{V}}^{+}$ $\mathcal{R}_{\mathsf{V}}^{-}$ 1-shot 2-shot 4-shot 8-shot 16-shot
[…]
3 ${}^{+}$ 62.83 63.31 63.95 65.13 66.27
4 ${}^{-}$ 62.65 63.07 63.60 64.36 65.12
5 DualAdapter 62.89 63.47 64.12 65.37 66.52
Effects of Different Visual Backbones. We also implement our DualAdapter with various visual encoders, including ResNet [17] and ViT [6], and evaluate its performance against other fine-tuning approaches in Fig. 6 (Left). We can see that our DualAdapter consistently exceeds other methods across all visual backbones, indicating the general effectiveness of our approach.
Figure 6: More ablation results. (Left) Performance comparison of our DualAdapter with other methods on few-shot learning task using different visual backbones; (Middle) Sensitivity analysis of $\lambda$ from Eq. (10) on ImageNet [4]; (Right) Sensitivity analysis of $\tau$ from Eq. (12) on 3 datasets, with the value of $\tau$ presented on a logarithmic scale.
[…]
Scalability to More Shots. When scaling up to more than 16 shots, Tip-Adapter-F [79] limits the cache size to prevent memory overflow. In contrast, our DualAdapter features reduced memory requirements, enabling it to operate without imposing cache size constraints, even in the 128-shot setting. Performance comparisons in Table 4.4 indicate that our DualAdapter outperforms Tip-Adapter-F [79] more significantly in higher shot settings, e.g., by 1.24% in the 128-shot setting.
Efficiency Comparison. We also compare the efficiency of DualAdapter with existing methods in Table 4.4. Our DualAdapter achieves the highest accuracy while also exhibiting advantageous computational efficiency: (1) Compared to prompt-based learning methods such as CoOp [84] and ProGrad [87], our DualAdapter requires approximately 300$\times$ less training time and over 1000$\times$ fewer FLOPs since we do not need to propagate gradients through the textual encoder; (2) Compared to adapter-style fine-tuning methods, our DualAdapter is also competitive in efficiency. It requires 10$\times$ less training time than CLIP-Adapter [12], and demands 3$\times$ fewer FLOPs and 4$\times$ fewer learnable parameters than Tip-Adapter-F [79].
Table 3: Performance comparison with Tip-Adapter(-F) [79] on ImageNet [4] with 16-/32-/64-/128-shot settings.
Table 4: Efficiency comparison with other existing methods on 16-shot ImageNet [4].
Method 16-shot 32-shot 64-shot 128-shot
CLIP [48] ${}^{\dagger}$0-shot: 60.33
Tip-Adapter [79] 62.03 62.51 62.88 63.15
Tip-Adapter-F [79] 65.47 66.58 67.96 69.74
DualAdapter (Ours) 66.52 67.68 69.01 70.98
Method Training Epochs GFLOPs Param. Accuracy
CLIP [48] - - - - 60.33
CoOp [84] 14 hr 200 >10 0.01M 62.95
ProGrad [87] 17 hr 200 >10 0.01M 63.45
CLIP-Adapter [12] 50 min 200 0.004 0.52M 63.59
Tip-Adapter-F [79] 5 min 20 0.030 16.38M 65.51
DualAdapter (Ours) 5 min 20 0.009 4.10M 66.52
Table 4: Efficiency comparison with other existing methods on 16-shot ImageNet [4].
5 Conclusion
In this work, we introduce DualAdapter, a novel approach to effectively adapt vision-language models for downstream tasks by integrating both positive and negative adapters. This innovative approach allows for effective few-shot adaptation from a dual perspective. Furthermore, we develop an unsupervised similarity-based label refinement approach to mitigate the adverse effects of low-quality image samples during the few-shot adaptation process. Comprehensive evaluations on 15 diverse datasets demonstrate that DualAdapter outperforms the state-of-the-art methods in both few-shot learning and domain generalization tasks.
Limitations. We identify two potential limitations of our DualAdapter: (1) Its efficacy in zero-shot situations is constrained due to the scarcity of original negative descriptions in the CLIP training corpus; (2) Similar to other adapter-style approaches such as Tip-Adapter-F [79], our DualAdapter fine-tuned on a specific task cannot be directly applied to another task without additional adaptation. However, recently, Wang et al. [68] shows that adapter-style fine-tuning methods can be extended for these scenarios using the $k$NN algorithm.
Acknowledgement
This work has been funded in part by the Army Research Laboratory (ARL) under grant W911NF-23-2-0007 and W911NF-19-2-0146, and the Air Force Office of Scientific Research (AFOSR) under grants FA9550-18-1-0097 and FA9550-18-1-0251.
Supplementary Material
In this supplementary document, we provide additional details and experimental results to enhance understanding and insights into our proposed DualAdapter. This supplementary document is organized as follows:
  • $\bullet$
    Full numerical results for the few-shot learning task are detailed in Section A.1.
  • $\bullet$
    We compare our DualAdapter with other state-of-the-art methods in domain generalization tasks, utilizing an alternative ViT backbone, in Section A.2.
  • $\bullet$
    More sensitivity analyses of the hyper-parameters are conducted in Section A.3.
  • $\bullet$
    Detailed statistics for all utilized datasets are provided in Section B.1.
  • $\bullet$
    We list the specific positive and negative prompts we used for each dataset in Section B.2.
  • $\bullet$
    Finally, we explore potential avenues for future research in Section C.
A Additional Experimental Results
A.1 Full Numerical Results on Few-Shot Learning
In Fig. 5 in Section 4.2 of the main text, we have evaluated our DualAdapter on few-shot learning task and compared with other state-of-the-art methods. In Table A1, we present the corresponding full numerical results on the few-shot learning task. We also report the 95% confidence interval over 3 random seeds of our DualAdapter to ensure reliability of our results. In the last column, we present the average recognition accuracy over 11 datasets. The results indicate that our DualAdapter consistently outperforms other state-of-the-art methods across various few-shot learning settings by substantial margins.
Our DualAdapter demonstrates superior recognition performance across nearly all tested scenarios, with certain exceptions in lower-shot settings for the Food101 [1] and OxfordPets [45] datasets. We attribute this to the prevalent challenge of overfitting, a common issue not exclusive to our approach but also affecting many existing methods, especially TaskRes [75]: While TaskRes consistently secures the second-best performance across the other 9 datasets, it underperforms significantly on these two. We hypothesize that this issue arises from the noisy training data with intense colors and sometimes wrong labels [1, 45]. However, in this work, we have designed a label refinement mechanism specifically to address this issue. As a result, our DualAdapter secures a relatively robust performance and achieves the second-best on these two datasets.
Table A1: Full numerical results on few-shot learning task. For each dataset, we report the mean accuracy and 95% confidence interval over 3 random seeds of our DualAdapter on 1-/2-/4-/8-/16-shot settings. ${}^{\dagger}$e report the zero-shot performance of CLIP [48] for all settings. For TaskRes [75], we report the results using the enhanced base classifier (i.e., TaskRes*). The best results are in bold and the second are underlined.
Method Setting
Caltech101 [11]
DTD [3]
EuroSAT [18]
FGVCAircraft [39]
Flowers102 [44]
Food101 [1]
ImageNet [4]
OxfordPets [45]
StanfordCars [29]
SUN397 [71]
UCF101 [57]
Avg.
${}^{\dagger}$ero-shot CLIP [48] 1-shot 84.52 40.33 41.80 16.98 65.46 77.31 60.33 85.51 54.26 58.56 61.44 58.77
CoOp [84] 87.43 44.13 50.51 9.80 67.90 73.71 57.15 86.51 55.48 60.10 62.10 59.53
CoCoOp [83] 86.01 45.14 35.08 17.81 67.52 77.42 60.84 86.96 57.22 62.28 62.84 59.92
CLIP-Adapter [12] 88.70 45.66 61.51 17.21 73.43 76.77 61.20 85.99 55.14 61.28 62.29 62.65
Tip-Adapter-F [79] 89.38 50.31 59.16 20.83 80.13 77.61 61.32 86.47 58.51 62.51 64.91 64.65
TaskRes [75] 88.80 50.20 61.70 21.41 79.17 74.03 61.90 83.60 59.13 62.33 64.77 64.28
90.87 53.13 67.70 23.61 84.17 77.93 62.89 86.90 61.34 65.13 68.25 67.45
DualAdapter (Ours) (±$\pm 0.33$) (±$\pm 0.48$) (±$\pm 0.47$) (±$\pm 0.18$) (±$\pm 0.34$) (±$\pm 0.22$) (±$\pm 0.13$) (±$\pm 0.31$) (±$\pm 0.28$) (±$\pm 0.08$) (±$\pm 0.23$) (±$\pm 0.28$)
${}^{\dagger}$ero-shot CLIP [48] 2-shot 84.52 40.33 41.80 16.98 65.46 77.31 60.33 85.51 54.26 58.56 61.44 58.77
CoOp [84] 87.92 45.04 60.43 18.25 77.47 72.26 55.88 82.36 58.10 59.82 64.13 61.97
CoCoOp [83] 89.47 46.20 38.51 20.22 70.70 78.81 61.86 88.81 58.28 63.50 65.23 61.96
CLIP-Adapter [12] 89.32 51.81 64.11 20.10 81.77 77.20 61.52 86.73 58.71 62.21 67.27 65.52
Tip-Adapter-F [79] 89.81 54.00 65.82 23.47 82.50 77.83 61.69 87.10 62.05 63.55 66.23 66.73
TaskRes [75] 90.27 55.13 65.83 24.13 86.57 75.17 62.63 84.63 63.70 64.97 70.00 67.54
91.19 59.17 74.40 26.22 88.43 78.35 63.47 87.68 64.77 66.73 70.25 70.06
DualAdapter (Ours) (±$\pm 0.26$) (±$\pm 0.31$) (±$\pm 0.87$) (±$\pm 0.37$) (±$\pm 0.18$) (±$\pm 0.17$) (±$\pm 0.06$) (±$\pm 0.28$) (±$\pm 0.21$) (±$\pm 0.33$) (±$\pm 0.48$) (±$\pm 0.31$)
${}^{\dagger}$ero-shot CLIP [48] 4-shot 84.52 40.33 41.80 16.98 65.46 77.31 60.33 85.51 54.26 58.56 61.44 58.77
CoOp [84] 89.17 53.38 70.20 21.72 85.81 72.72 59.93 87.22 61.92 63.46 67.08 66.60
CoCoOp [83] 90.31 47.90 63.56 20.56 72.72 79.51 62.52 88.60 59.90 64.90 67.90 65.31
CLIP-Adapter [12] 89.98 57.02 73.18 22.99 87.30 77.93 61.84 87.36 62.26 65.90 68.90 68.61
Tip-Adapter-F [79] 90.67 57.78 73.85 26.01 89.02 78.26 62.52 87.72 64.82 66.13 70.87 69.79
TaskRes [75] 90.97 60.70 73.83 25.70 90.20 76.10 63.57 86.33 67.43 67.27 70.93 70.28
92.21 66.01 76.54 28.95 92.04 78.74 64.12 88.13 67.96 68.59 73.46 72.43
DualAdapter (Ours) (±$\pm 0.21$) (±$\pm 0.49$) (±$\pm 0.59$) (±$\pm 0.29$) (±$\pm 0.25$) (±$\pm 0.11$) (±$\pm 0.17$) (±$\pm 0.26$) (±$\pm 0.36$) (±$\pm 0.29$) (±$\pm 0.51$) (±$\pm 0.32$)
${}^{\dagger}$ero-shot CLIP [48] 8-shot 84.52 40.33 41.80 16.98 65.46 77.31 60.33 85.51 54.26 58.56 61.44 58.77
CoOp [84] 90.15 59.88 76.51 25.93 90.84 71.52 60.91 86.40 68.49 65.63 71.81 69.82
CoCoOp [83] 90.14 52.21 64.13 22.03 75.88 79.59 62.40 88.74 60.87 65.37 68.25 66.33
CLIP-Adapter [12] 91.22 60.70 78.34 25.77 91.79 78.01 62.68 87.70 67.78 67.52 73.02 71.32
Tip-Adapter-F [79] 91.54 62.67 77.83 30.21 91.85 78.71 64.00 88.07 69.53 68.80 74.50 72.52
TaskRes [75] 92.40 64.77 79.33 31.48 94.73 76.40 64.67 87.17 71.83 68.73 75.33 73.35
93.40 67.78 81.62 33.90 95.23 79.23 65.37 89.29 72.08 70.93 76.84 75.06
DualAdapter (Ours) (±$\pm 0.23$) (±$\pm 0.55$) (±$\pm 0.54$) (±$\pm 0.63$) (±$\pm 0.26$) (±$\pm 0.21$) (±$\pm 0.09$) (±$\pm 0.21$) (±$\pm 0.51$) (±$\pm 0.15$) (±$\pm 0.39$) (±$\pm 0.34$)
${}^{\dagger}$ero-shot CLIP [48] 16-shot 84.52 40.33 41.80 16.98 65.46 77.31 60.33 85.51 54.26 58.56 61.44 58.77
CoOp [84] 91.61 63.11 82.36 31.01 94.39 73.80 62.95 87.30 72.51 69.11 75.70 73.07
CoCoOp [83] 90.90 57.53 70.77 22.40 79.14 79.68 62.71 89.93 62.22 67.21 70.81 68.48
CLIP-Adapter [12] 92.44 66.14 82.76 31.83 93.91 78.21 63.59 87.91 74.12 69.59 76.80 74.30
Tip-Adapter-F [79] 92.93 67.33 83.80 35.50 95.01 79.50 65.51 89.71 75.50 71.31 78.01 75.83
TaskRes [75] 93.43 67.13 84.03 36.30 96.03 77.60 65.73 87.83 76.83 70.67 77.97 75.78
93.77 70.83 87.36 40.27 96.51 79.87 66.52 90.58 77.48 72.32 80.28 77.80
DualAdapter (Ours) (±$\pm 0.33$) (±$\pm 0.77$) (±$\pm 0.84$) (±$\pm 0.53$) (±$\pm 0.58$) (±$\pm 0.34$) (±$\pm 0.13$) (±$\pm 0.38$) (±$\pm 0.68$) (±$\pm 0.24$) (±$\pm 0.26$) (±$\pm 0.50$)
A.2 More Results on Domain Generalization
Table A2: Performance comparison on robustness to distribution shifts. All the models are trained on 16-shot ImageNet [4] and directed tested on the OOD target datasets. The best results are in bold and the second best are underlined.
Method Source Target
ImageNet -V2 -Sketch -A -R Avg.
Zero-Shot CLIP [48] 62.05 54.79 40.82 29.57 65.99 47.79
Linear Probe CLIP [48] 59.58 49.73 28.06 19.67 47.20 36.17
CoOp [84] 66.85 58.08 40.44 30.62 64.45 48.40
TaskRes [75] 68.20 59.20 42.50 31.43 69.33 50.62
GraphAdapter [32] 68.47 59.10 42.70 31.73 69.43 50.74
DualAdapter (Ours) 69.63 59.76 43.41 32.48 69.60 51.31
In Table 1 in Section 4.3 of the main text, we compare the generalization performance of our DualAdapter with other state-of-the-art methods in the presence of distribution shifts with ResNet-50 visual backbone. In Table A2, we present the performance comparison on domain generalization task using ViT-B/32 visual backbone. Consistently, our DualAdapter not only achieves state-of-the-art performance on the source dataset but also attains an average performance gain of 0.57% across 4 out-of-distribution (OOD) target datasets. This verifies that our DualAdapter demonstrates superior generalizability compared to other state-of-the-art methods, independent of the visual backbone utilized.
A.3 More Sensitivity Analyses of Hyper-Parameters
Table A3: Sensitivity of hyper-parameters. All the results are reported on 16-shot ImageNet [4].
$\alpha$ 0.0 0.5 1.0 1.2 1.5 2.0 66.34 66.40 66.47 66.52 66.44 66.36 $\beta$ 1.0 1.5 2.0 2.5 3.0 3.5 66.38 66.44 66.52 66.50 66.48 66.40
Building upon the sensitivity analyses of $\lambda$ and $\tau$ detailed in Section 4.4 of the main text, this section extends our examination to include the sensitivity of parameters $\alpha$ and $\beta$ on 16-shot ImageNet [4]. In our experiments on ImageNet [4], we set the hyper-parameters $\alpha$ and $\beta$ defined in Section 3 to $1.2$ and $2.0$, respectively. To comprehensively investigate the effects of different hyper-parameters, we conducted a sensitivity experiment where we varied each hyper-parameter individually and evaluated the performance on 16-shot ImageNet [4] in Fig. A3. We can see that our choice of $\alpha=1.2$ and $\beta=2.0$ yields the highest performance. Moreover, our DualAdapter maintains robust performance when adjusting these two hyper-parameters, since our DualAdapter includes adapters from both textual and visual modalities and each of our four adapters can work effectively, as we presented in Table 2 in the main text.
B Additional Implementation Details
B.1 Dataset Details
In Table B4, we present the detailed statistics of each dataset we used in our experiments, including the number of classes, the sizes of training, validation and testing sets, and their original tasks.
Table B4: Detailed statistics of datasets used in experiments. Note that the last 4 ImageNet variant datasets are designed for evaluation and only contain the test sets.
Dataset Classes Training Validation Testing Task
Caltech101 [11] 100 4,128 1,649 2,465 Object recognition
DTD [3] 47 2,820 1,128 1,692 Texture recognition
EuroSAT [18] 10 13,500 5,400 8,100 Satellite image recognition
FGVCAircraft [39] 100 3,334 3,333 3,333 Fine-grained aircraft recognition
Flowers102 [44] 102 4,093 1,633 2,463 Fine-grained flowers recognition
Food101 [1] 101 50,500 20,200 30,300 Fine-grained food recognition
ImageNet [4] 1,000 1.28M - 50,000 Object recognition
OxfordPets [45] 37 2,944 736 3,669 Fine-grained pets recognition
StanfordCars [29] 196 6,509 1,635 8,041 Fine-grained car recognition
SUN397 [71] 397 15,880 3,970 19,850 Scene recognition
UCF101 [57] 101 7,639 1,898 3,783 Action recognition
ImageNet-V2 [50] 1,000 - - 10,000 Robustness of collocation
ImageNet-Sketch [65] 1,000 - - 50,889 Robustness of sketch domain
ImageNet-A [20] 200 - - 7,500 Robustness of adversarial attack
ImageNet-R [19] 200 - - 30,000 Robustness of multi-domains
B.2 Positive and Negative Prompts
In Table B5, we detail the specific positive and negative prompts utilized for each dataset. Additionally, as mentioned in Section 4.1, we incorporate prompts from CuPL [46] to further enhance model performance.
Table B5: Positive and negative prompts used in experiments. In addition to these prompts, we also employ CuPL [46] prompts to further enhance performance.
Dataset Positive Prompts Negative Prompts
“itap of a {CLASS}.” “itap without any {CLASS}.”
ImageNet [4] “a bad photo of the {CLASS}.” “a bad photo with no {CLASS} in it.”
ImageNet-V2 [50] “a origami {CLASS}.” “a origami that isn’t a {CLASS}.”
ImageNet-Sketch [65] “a photo of the large {CLASS}.” “a photo with no large {CLASS}.”
ImageNet-A [20] “a {CLASS} in a video game.” “a video game scene without a {CLASS}.”
ImageNet-R [19] “art of the {CLASS}.” “art that doesn’t include a {CLASS}.”
“a photo of the small {CLASS}.” “a photo with no small {CLASS}.”
Caltech101 [11] “a photo of a {CLASS}.” “a photo without {CLASS}.”
DTD [3] “{CLASS} texture.” “not {CLASS} texture.”
EuroSAT [18] “a centered satellite photo of {CLASS}.” “a centered satellite photo without {CLASS}.”
FGVCAircraft [39] “a photo of a {CLASS}, a type of aircraft.” “a photo without {CLASS}, a type of aircraft.”
Flowers102 [44] “a photo of a {CLASS}, a type of flower.” “a photo without {CLASS}, a type of flower.”
Food101 [1] “a photo of {CLASS}, a type of food.” “a photo without {CLASS}, a type of food.”
OxfordPets [45] “a photo of a {CLASS}, a type of pet.” “a photo without {CLASS}, a type of pet.”
StanfordCars [29] “a photo of a {CLASS}.” “a photo of no {CLASS}.”
SUN397 [71] “a photo of a {CLASS}.” “a photo without {CLASS}.”
UCF101 [57] “a photo of a person doing {CLASS}.” “a photo of a person not doing {CLASS}.”
C Future Work
In this work, we introduce the first dual-path adapter-style fine-tuning framework for VLMs. We believe that the similar concept of dual learning can also be applied to prompt-based learning methods, and be extended to fine-tune other foundational models (e.g., other VLMs [24, 74] and LLMs [63, 5]). Besides, we also notice that there are a lot of research efforts dedicated to design better prompts (e.g., using LLM [46, 40, 70]) to fully exploit the capabilities of CLIP. We hope that with our work, future research endeavors can also be directed to investigate the utilization of negative prompts to better activate the negative inference capabilities, further broadening the scope and effectiveness of CLIP.